Logiweb dictionary of problemtwo

Up Help

0 0 problemtwo

1 0 L_{o} 2 0 L_{o} 3 0 L_{o} 4 0 L_{o} 5 0 L_{o} 6 0 L_{o} 7 0 L_{o} 8 0 L_{o} 9 0 L_{o} 10 0 L_{o} 11 0 L_{o} 12 0 L_{o} 13 0 L_{o} 14 0 L_{o} 15 0 L_{o} 16 0 L_{o} 17 0 L_{o} 18 0 L_{o} 19 0 L_{o}

20 0 pred calc
21 0 pc1
22 0 pc2
23 0 pc3
24 0 pc4
25 0 pc5
26 0 pc6
27 0 pc7
28 0 pc8

29 0 pc9
30 0 pc10
31 0 pc11
32 0 pc12
33 0 pcmp
34 0 pcunsound
35 0 pcded
36 0 pcia
37 0 pcie
38 0 pcdeduction
39 0 trivia
40 0 trivia2
41 0 iatest
42 0 andintro
43 0 andelim1
44 0 andelim2
45 0 orintro1
46 0 orintro2
47 0 orelim
48 0 notintro
49 0 notnotintro
50 0 notnotelim
51 0 mt
52 0 pbc
53 0 repeat
54 0 lem
55 0 forallintro
56 0 forallelim
57 0 existsintro
58 0 existselim
59 0 bottomelim
60 0 lemnotintro
61 0 nontriv0
62 0 nontriv1

63 0 nontriv2
64 0 nontriv3
65 0 nontriv4
66 0 nontriv5
67 0 nontriv6

68 2 ∗ ≡ ∗ 69 2 ∗ = ∗ 70 1 ¬∗ 71 2 ∗ ∧ ∗ 72 2 ∗ ∨ ∗ 73 2 ∀ ∗ . (∗) 74 2 ∃ ∗ . (∗) 75 2 ∗ ∈ ∗

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-07-12.UTC:19:00:11.024866 = MJD-53928.TAI:19:00:44.024866 = LGT-4659447644024866e-6